# Machine learning-based identification of Gaia astrometric exoplanet orbits

### v1

@@ Abstract
The third Gaia data release (DR3) contains $\sim$170 000 astrometric orbit solutions of two-body systems located within $\sim$500 pc of the Sun. Determining component masses in these systems, in particular of stars hosting exoplanets, usually hinges on incorporating complementary observations in addition to the astrometry, e.g. spectroscopy and radial velocities. Several DR3 two-body systems with exoplanet, brown-dwarf, stellar, and black-hole components have been confirmed in this way. We developed an alternative machine learning approach that uses only the DR3 orbital solutions with the aim of identifying the best candidates for exoplanets and brown-dwarf companions. Based on confirmed substellar companions in the literature, we use semi-supervised anomaly detection methods in combination with extreme gradient boosting and random forest classifiers to determine likely low-mass outliers in the population of non-single sources. We employ and study feature importance to investigate the method’s plausibility and produced a list of 22 best candidates of which four are exoplanet candidates and another five are either very-massive brown dwarfs or very-low mass stars. Three candidates, including one initial exoplanet candidate, correspond to false-positive solutions where longer-period binary star motion was fitted with a biased shorter-period orbit. We highlight nine candidates with brown-dwarf companions for preferential follow-up. One candidate companion around the Sun-like star G 15-6 could be confirmed as a genuine brown dwarf using external radial-velocity data. This new approach is a powerful complement to the traditional identification methods for substellar companions among Gaia astrometric orbits. It is particularly relevant in the context of Gaia DR4 and its expected exoplanet discovery yield.
@@ 1 Introduction
The third Gaia data release (DR3, Gaia Collaboration et al., 2016, 2023a) delivered the first uniform and large-scale census of astrometric binaries and increased the number of available orbital solutions by two orders of magnitude from a few thousand (Mason et al., 2001) to $\sim$170 000. The revolutionary accuracy of Gaia made it possible to determine a large number of astrometric orbits with semi-major axes smaller than 1 milli-arcsecond (mas) for the first time (Halbwachs et al., 2023), finally opening the astrometric exoplanet-science window widely. However, small semi-major axes do not necessarily imply low companion masses: because the DR3 astrometric orbits usually refer to the unresolved photocentre of a binary, the principal astrophysical false-positive scenario for an exoplanet candidate is a nearly equal-mass and equal-brightness binary star (e.g. Gaia Collaboration et al., 2023b; Marcussen & Albrecht, 2023).
The astrometric orbits of several known giant exoplanets have been determined in DR3 (Holl et al., 2023; Gaia Collaboration et al., 2023b; Winn, 2022), leading to better constraints on their orbital configurations and masses. Gaia DR3 also contains a number of new giant exoplanet candidates from astrometry. [^1] The Gaia exoplanet candidate list is at https://www.cosmos.esa.int/web/gaia/exoplanets The default procedure to confirm those is to gather additional data, e.g. spectroscopy and precision radial-velocities (RV), that exclude any false-positive scenario.
Here we propose an intermediate step to identify the most promising exoplanet candidates that can be prioritised for follow-up. Our machine learning approach relies only on DR3 without the need for external datasets. Because of the very small number of confirmed exoplanets with DR3 orbits, which we need to ‘train‘ our models, we use as proof-of-concept the identification of substellar companions in general, i.e. both extrasolar planets and brown-dwarf companions. Since the mass distribution of substellar companions is structured but continuous (e.g. Grether & Lineweaver, 2006; Sahlmann et al., 2011), this does not imply a loss of applicability, but it conveniently doubles the size of the ‘training‘ sample.
Figure 1: Overview of the data and different partitionings used. Source preselection was performed as described in Section 2.2.
@@ 2 Gaia DR3 Data selection and augmentation
We queried the gaiadr3.nss_two_body_orbit table and selected all astrometric orbits, i.e. entries having a nss_solution_type of either ‘Orbital’, ‘AstroSpectroSB1’, ‘OrbitalAlternative*’, or ‘OrbitalTargetedSearch*’.
For 98 sources that have both ‘AstroSpectroSB1’ and ‘OrbitalTargetedSearch*’ solutions we only retained the ‘AstroSpectroSB1’ solutions. Two solutions were identified as false positives by the Gaia collaboration [^2] https://www.cosmos.esa.int/web/gaia/dr3-known-issues#FalsePositive and we discarded those. This left us with 169 127 orbit solutions for the same number of sources.
For these sources, we retrieved the data of the gaia_source table which contains magnitudes, colours, radial velocities, and other information obtained from the astrometric single-star solution (Lindegren et al., 2021a).
The DR3 table fields are described in the online documentation (van Leeuwen et al., 2022), but not all are relevant for the purpose of our study. The selection of relevant fields and the augmentation with additional fields is described below.
The gaia_source and nss_two_body_orbit fields that we used are listed in Tables 5 and 6, respectively. We augmented these tables with a few fields to make relevant source or orbit parameters directly accessible and interpretable. For the gaia_source table, we computed the absolute source $G$-band magnitude from the apparent magnitude phot_g_mean_mag and the parallax $\varpi_{\mathrm{ss}}$ from gaia_source without zero-point correction (Lindegren et al., 2021b).
@@
where $P$ is the orbital period in days and $\varpi_{\mathrm{nss}}$ is the parallax from nss_two_body_orbit in mas. Finally we joined both tables on the source identifier source_id.
We note that the nss_solution_type is disregarded in the machine learning analysis. This is intentional because we want to apply our approach to purely astrometric orbits with as little selection effects as possible. A nss_solution_type = ‘AstroSpectroSB1’, for instance, implies that a compatible orbit was detected in two Gaia instruments at the same time, and the necessary RV amplitude makes it unlikely that an exoplanet signature corresponds to a ‘AstroSpectroSB1’.
@@ 2.1 Data imputation
For some sources/solutions not all table fields have been filled, e.g. sources that are too faint to be observed with the Gaia Radial Velocity Spectrometer (RVS, Katz et al., 2023) do not have an entry in the radial_velocity field. We therefore filled these occurrences with appropriate numerical values. The affected fields and fill-values are listed in Table 1. The fill values were determined as the median of the remaining population, expect for radial_velocity and eccentricity_error where they were set to zero.
@@ 2.2 Definition of labels
131 142 of our selected sources have entries in binary_masses. 105 of those have several mass estimates stemming with different binary_masses.combination_method. We removed those listed as ‘SB2’ and ‘Eclipsing’ which left us with 131 037 unique sources. The preselected sources label was assigned to 1838 sources with binary_masses.m2_lower $<0.08$ and the high-probability preselected sources label was assigned to sources with binary_masses.m2_upper $<0.08$ (cf. Stevenson et al., 2023). In the context, the preselected sources label refers to solutions that correspond to sources, which based on their entries in binary_masses could be substellar-companion candidates.
@@ 3 Data analysis
Figure 2: Overview of the algorithm used for identifying candidates.
3.1 Substellar Companions as Anomalies
Fundamentally, we expect the number of substellar companions in the non-single sources to be quite small, in the range of less than one percent and in the single digit percentile range amongst the preselected sources. As such, we can consider them to be outliers in our data, as can be appreciated in Figure 3, and we expect them to be identifiable based on their DR3 parameters.
Even though some relations between the astrometric data, such as the value of the mass function $f_{M}$ (see Equation 1), can be strongly indicative of the potential presence of a substellar companion, it is conceivable that there are various correlations and interactions between different parameters and the likelihood of a specific source having a substellar companion can likely not easily be reduced to a single parameter. Conversely, there are likely also other forms of outliers in the data, such as anomalies in the data processing pipeline or very low-mass stellar companions. Thus, we formulate the problem of identifying likely exoplanets and brown-dwarf companions as two-fold: first, in identifying outliers in the dataset and, secondly, identifying those outliers that are particularly likely to be substellar companions.
@@ 3.2 Semi-supervised Outlier Detection
In most established machine learning approaches (Zhao & Hryniewicki, 2018; Han et al., 2022), the task of semi-supervised outlier detection is approached by training on the limited amount of data containing both known outliers and inliers and then predicting on a separate dataset. Since we only have a few confirmed outliers and virtually no confirmed inliers (i.e. binaries and very low mass stellar companions) we developed a different approach and instead consider our entire dataset of inliers (i.e. the vast majority of the data) to be weakly labelled, i.e. potentially mislabelled. Thus, we transform the task from classifying newly unseen samples into inliers or outliers into identifying the likeliest mislabelled samples in our training data. To this end, we perform a k-fold cross-validation and use the predictions on the presumed inliers as the indication of whether we believe that sample to be a mislabelled one. This procedure is shown in detail in Figure 2. We choose the amount of folds to be identical to the number of outliers to maximise the amount of confirmed outliers in the training data. The left out outlier is used to compute the validation metrics in the results section.
To further improve the robustness of the approach and since precise performance evaluation is challenging with the limited amount of data , we use two different approaches in our ensembles. In particular, we utilise XGBoost (Chen & Guestrin, 2016) as implemented in the Python package of the same name [^3] https://xgboost.readthedocs.io/en/stable/index.html Accessed: 2024-03-08 and a balanced random forest classifier (Chen et al., 2004) as implemented in the imblearn [^4] https://imbalanced-learn.org/ Accessed: 2024-03-08 Python module. Given the invariance of both approaches to monotonous transforms, no data transforms were applied beyond the elimination of samples (i.e. sources/orbital solutions) missing any of the studied features.
@@ 3.3 Feature Selection and Importance
As examples, we discarded both the parallax and the apparent magnitude of the source. These correlate positively with having confirmed substellar companions because those are the best-surveyed sources with ground-based, high-precision radial-velocity observations and generally are easier to follow-up. From an statistical astrophysical standpoint, there is little reason for a bright source to be more likely to host a substellar companion than a faint one, at least within the parameter space surveyed here. The other confounders that we identified are indicated in Tables 5 and 6.
@@ 3.4 Ensembling and Candidate Selection
As outlined in the previous paragraphs, we use four different datasets (see Figure 1) and two different models (XGBoost and BalancedRandomForest). From these, we build our predictions using max-voting on the 50 samples with the highest outlier prediction as computed using the algorithm illustrated in Figure 2. We weight the voting relative to the chance to occur. Individual samples may, respectively, occur as inliers zero (confirmed substellar companions), two (high-probability preselected sources, orange in Figure 1), four (non-single sources) or eight (preselected sources) times. We measure the actual occurrence with three values: $\rho$ which describes the rate of occurrence in the eight runs and $\rho_{nss}$ and $\rho_{ssc}$, which are the rates of occurrence in the large datasets (D1 & D2) and smaller ones (D3 & D4), respectively.
@@ 4 Results: substellar-companion candidates
Refer to captionRefer to caption
Figure 3: Density histograms of absolute magnitude as a function of colour (top panel, the inset show a zoom into the region of interest) and mass function (bottom) for all DR3 astrometric orbits. The blue contours indicate the concentration of the solutions labelled as preselected sources. Circles indicate the 10 confirmed exoplanets, squares indicate the 14 confirmed BD-companions, and crosses indicate our 22 best candidates for substellar companions discussed in the text. For reference, the Sun has $M_{G}=4.67$ and $G_{\mathrm{BP}}-G_{\mathrm{RP}}=0.82$ (Casagrande & VandenBerg, 2018), and the mass function of a $5\,M_{J}$ planet in a Jupiter-like orbit around the Sun is $1.1\cdot 10^{-7}$. These reference locations are marked with a yellow diamond. The exoplanet with the largest mass function is HD 39392 b (Wilson et al., 2016; Stevenson et al., 2023).
@@ 4.1 Validation
Ensuring the performance of our approach is a challenging task given the small amount of labelled samples. In the prediction method described in Figure 2 we obtain the predictions on the left-out fold as in a stratified K-fold validation. We use these predictions to compute performance metrics. As the inlier labels in these folds are weak labels (we expect some substellar companions in them) the metrics are probably underestimating the true performance since the models correctly "mislabel" some false inliers. Similarly, for the high probability (but not confirmed) outliers some may be mislabelled outliers, further impacting the performance measurements. Detailed results for different configurations are given in Table 3. Overall, we achieve very good results on the area under the receiver operating characteristic (AUROC) and accuracy of predictions. The area under the precision recall curve (AUPRC) results are worse but this is expected given the weak labels and significant class imbalance (less than 0.03% labelled outliers amongst all sources).
Table 3: Validation results for different configurations showing accuracy of prediction as well as area under the receiver operating characteristic (AUROC) and area under the precision recall curve (AUPRC).
Metric Only Confirmed Outliers High Probability Outliers All Sources Candidate Sources XGBoost Random Forest All
@@
A secondary measure of the performance of our approach is the accuracy of the applied majority voting scheme to identify the best candidates from the eight configurations. A total of 170 unique sources show up in the top 50 outliers of the eight approaches (i.e. amongst all samples including presumed inliers and outliers). Among those 170, we find 22 of our 24 confirmed substellar companions among the 61 highest ranked samples across all configurations, each being identified in at least two separate runs. The only missing ones are the brown dwarf companions in LHS 1610 (Gaia DR3 43574131143039104, Fitzmaurice et al., 2023) and the brown dwarf in HD 89707 (Gaia DR3 3751763647996317056, Gaia Collaboration et al., 2023b). The one around LHS 1610 was likely missed due to the very-red colour of its M5-dwarf host star, which is uncommon in the confirmed substellar companions. The one around HD 89707 was barely below the detection threshold scoring highly on most models. We investigate this further in Section 4.4 on feature importance.
Table 4: Best candidates for exoplanets and brown-dwarf companions. The SpT column indicates the source’s spectral type as listed in Simbad, the solution type column indicates a short-hand for the nss_solution_type (similar to Marcussen & Albrecht 2023: OTSV is ‘OrbitalTargetedSearchValidated’, OTS is ‘OrbitalTargetedSearch’, and ASB1 is ‘AstroSpectroSB1’), $M_{2}$ is a companion-mass estimate under the dark-companion assumption and using the primary-mass estimate $M_{1}$ from binary_masses, $\rho$ is the relative occurrence introduced in Section 3.4, $\rho_{\mathrm{nss}}$ is the relative occurrence for dataset variants D1&D2 only, $\rho_{\mathrm{ssc}}$ is the relative occurrence for dataset variants D3&D4 only, the $ssc$ flag is 1 when the solution was part of ‘preselected sources ‘ or ‘high-probability preselected sources ‘ and 0 otherwise, and $M_{2,\mathrm{alt}}$ is the companion-mass range as estimated in binary_masses.

### v2

@@ Abstract
The third Gaia data release (DR3) contains $\sim$170 000 astrometric orbit solutions of two-body systems located within $\sim$500 pc of the Sun. Determining component masses in these systems, in particular of stars hosting exoplanets, usually hinges on incorporating complementary observations in addition to the astrometry, e.g. spectroscopy and radial velocities. Several Gaia DR3 two-body systems with exoplanet, brown-dwarf, stellar, and black-hole components have been confirmed in this way. We developed an alternative machine learning approach that uses only the Gaia DR3 orbital solutions with the aim of identifying the best candidates for exoplanets and brown-dwarf companions. Based on confirmed substellar companions in the literature, we use semi-supervised anomaly detection methods in combination with extreme gradient boosting and random forest classifiers to determine likely low-mass outliers in the population of non-single sources. We employ and study feature importance to investigate the method’s plausibility and produced a list of 20 best candidates of which two are exoplanet candidates and another five are either very-massive brown dwarfs or very-low mass stars. Three candidates, including one initial exoplanet candidate, correspond to false-positive solutions where longer-period binary star motion was fitted with a biased shorter-period orbit. We highlight nine candidates with brown-dwarf companions for preferential follow-up. The companion around the Sun-like star G 15-6 could be confirmed as a genuine brown dwarf using external radial-velocity data. This new approach is a powerful complement to the traditional identification methods for substellar companions among Gaia astrometric orbits. It is particularly relevant in the context of Gaia DR4 and its expected exoplanet discovery yield.
@@ 1 Introduction
The third Gaia data release (DR3, Gaia Collaboration et al., 2016, 2023a) delivered the first uniform and large-scale census of astrometric binaries and increased the number of available orbital solutions by two orders of magnitude from a few thousand (Mason et al., 2001) to $\sim$170 000. The revolutionary accuracy of Gaia made it possible to determine a large number of astrometric orbits with semi-major axes smaller than 1 milli-arcsecond (mas) for the first time (Halbwachs et al., 2023), finally opening the astrometric exoplanet-science window widely. However, small semi-major axes do not necessarily imply low companion masses: because the Gaia DR3 astrometric orbits usually refer to the unresolved photocentre of a binary, the principal astrophysical false-positive scenario for an exoplanet candidate is a nearly equal-mass and equal-brightness binary star (e.g. Gaia Collaboration et al., 2023b; Marcussen & Albrecht, 2023).
The astrometric orbits of several known giant exoplanets have been determined in Gaia DR3 (Holl et al., 2023; Gaia Collaboration et al., 2023b; Winn, 2022), leading to better constraints on their orbital configurations and masses. Gaia DR3 also contains a number of new giant exoplanet candidates from astrometry. [^1] The Gaia DR3 exoplanet candidate list is at https://www.cosmos.esa.int/web/gaia/exoplanets The default procedure to confirm those is to gather additional data, e.g. spectroscopy and precision radial-velocities (RV), that exclude any false-positive scenario.
Here we propose an intermediate step to identify the most promising exoplanet candidates that can be prioritised for follow-up. Our machine learning approach relies only on Gaia DR3 without the need for external datasets. Because of the very small number of confirmed exoplanets with Gaia DR3 orbits, which we need to ‘train‘ our models, we use as proof-of-concept the identification of substellar companions in general, i.e. both extrasolar planets and brown-dwarf companions. Since the mass distribution of substellar companions is structured but continuous (e.g. Grether & Lineweaver, 2006; Sahlmann et al., 2011), this does not imply a loss of applicability, but it conveniently doubles the size of the ‘training‘ sample.
Figure 1: Overview of the data and different partitionings (Datasets D1, D2, D3, D4) used. Source preselection was performed as described in Section 2.2.
@@ 2 Gaia DR3 Data selection and augmentation
The principle data source for this study are two tables in the Gaia DR3 archive [^2] https://gea.esac.esa.int/archive/: first the main astrometric catalog published in a table named gaiadr3.gaia_source resulting from fitting a single-star model, and second the Keplerian-orbit solutions recorded in gaiadr3.nss_two_body_orbit, which were obtained by fitting a non-single star (NSS) model. We queried the nss_two_body_orbit table and selected all astrometric orbits, i.e. entries having a nss_solution_type of either ‘Orbital’, ‘AstroSpectroSB1’, ‘OrbitalAlternative*’, or ‘OrbitalTargetedSearch*’ 333‘OrbitalAlternative*’ is a shorthand for either ‘OrbitalAlternative’ or ‘OrbitalAlternativeValidated’ and ‘OrbitalTargetedSearch*’ is a shorthand for ‘OrbitalTargetedSearch’ or ‘OrbitalTargetedSearchValidated’.. The former two types were produced by the ‘binary pipeline‘ (Halbwachs et al., 2023), whereas ‘OrbitalAlternative*’ and ‘OrbitalTargetedSearch*’ solutions were produced by the ‘exoplanet pipeline‘ (Holl et al., 2023). The exoplanet pipeline produced 198 solutions of type ‘OrbitalTargetedSearchValidated’, which correspond to systems whose orbits where confirmed with independent datasets (Holl et al., 2023).
For 98 sources that have both ‘AstroSpectroSB1’ and ‘OrbitalTargetedSearch*’ solutions we only retained the ‘AstroSpectroSB1’ solutions. Two solutions were initially identified as false positives by the Gaia collaboration [^4] https://www.cosmos.esa.int/web/gaia/dr3-known-issues#FalsePositive and we discarded those. Two more false-positive solutions were reported during the refereeing process, which we did not account for here but they are discussed in Section 4.2. This left us with 169 127 orbit solutions for the same number of sources.
For these sources, we retrieved the data of the gaia_source table which contains magnitudes, colours, radial velocities, and other information obtained from the astrometric single-star solution (Lindegren et al., 2021a). The Gaia DR3 table fields are described in the online documentation (van Leeuwen et al., 2022), but not all are relevant for the purpose of our study. The selection of relevant fields and the augmentation with additional fields is described below.
The gaia_source and nss_two_body_orbit fields that we used are listed in Tables 5 and 6, respectively. We augmented these tables with a few fields to make relevant source or orbit parameters directly accessible and interpretable. For the gaia_source table, we computed the absolute source $G$-band magnitude from the apparent magnitude phot_g_mean_mag and the single-star model parallax $\varpi_{\mathrm{ss}}$ from gaia_source without zero-point correction (Lindegren et al., 2021b).
@@
where $P$ is the orbital period in days and $\varpi_{\mathrm{nss}}$ is the non-single-star model parallax from nss_two_body_orbit in mas. Finally we joined both tables on the source identifier source_id.
We note that the nss_solution_type is disregarded in the machine learning analysis. This is intentional because we want to apply our approach to purely astrometric orbits with as little selection effects as possible. A nss_solution_type = ‘AstroSpectroSB1’, for instance, implies that a compatible orbit was detected in two Gaia instruments at the same time, imposing additional constraints on source apparent magnitude, and the necessary RV amplitude makes it unlikely that an exoplanet signature corresponds to a ‘AstroSpectroSB1’.
@@ 2.1 Data imputation
For some sources/solutions not all table fields have been filled, e.g. sources that are too faint to be observed with the Gaia Radial Velocity Spectrometer (RVS, Katz et al., 2023) do not have an entry in the radial_velocity field. We therefore filled these occurrences with appropriate numerical values. The affected fields and fill-values are listed in Table 1. The fill values were determined as the median of the remaining population, except for radial_velocity and eccentricity_error where they were set to zero [^5] The median radial_velocity of sources where it is available is $-2.48$ km/s, thus generally the sample average value is compatible with zero..
@@ 2.2 Definition of labels
131 142 of our selected sources have entries in binary_masses. 105 of those have several mass estimates originating from different mass estimation methods as indicated by the field binary_masses.combination_method. We removed entries listed with a combination method that includes ‘SB2’ and ‘Eclipsing’, i.e. that rely on a double-lined spectroscopic or eclipsing binary solution, respectively, which left us with 131 037 unique sources. The preselected sources label was assigned to 1838 sources with binary_masses.m2_lower $<0.08$ and the high-probability preselected sources label was assigned to sources with binary_masses.m2_upper $<0.08$ (cf. Stevenson et al., 2023). In the context, the preselected sources label refers to solutions that correspond to sources, which based on their entries in binary_masses could be substellar-companion candidates.
@@ 3 Data analysis
Figure 2: Overview of the system architecture used for identifying candidates.
3.1 Substellar Companions as Outliers
Fundamentally, we expect the number of substellar companions in the non-single sources to be quite small, in the range of less than one percent and in the single digit percentile range amongst the preselected sources. As such, we can consider them to be outliers in our data and we expect them to be identifiable based on their Gaia DR3 parameters.
Even though some relations between the astrometric data, such as the value of the mass function $f_{M}$ (see Equation 1), can be strongly indicative of the potential presence of a substellar companion, it is conceivable that there are various correlations and interactions between different parameters and the likelihood of a specific source having a substellar companion can likely not easily be reduced to a single parameter. Conversely, there are likely also other forms of outliers in the data, such as anomalies in the data processing pipeline or very low-mass stellar companions. Thus, we formulate the problem of identifying likely exoplanets and brown-dwarf companions as two-fold: first, in identifying outliers in the dataset and, secondly, identifying those outliers that are particularly likely to be substellar companions. We achieve this by formulating a semi-supervised outlier detection problem, in which the outliers used in training are previously identified substellar companions and not other outliers such as data processing anomalies or very low-mass stellar companions.
@@ 3.2 Semi-supervised Outlier Detection
In most established machine learning approaches (Zhao & Hryniewicki, 2018; Han et al., 2022), the task of semi-supervised outlier detection is approached by training on the limited amount of data containing both known outliers and inliers and then predicting on a separate dataset. Since we only have a few confirmed outliers and virtually no confirmed inliers (i.e. binaries and very low mass stellar companions) we developed a different approach and instead consider our entire dataset of inliers (i.e. the vast majority of the data) to be weakly labelled, i.e. potentially mislabelled. Thus, we transform the task from classifying newly unseen samples into inliers or outliers into identifying the likeliest mislabelled samples in our training data. To this end, we perform a k-fold cross-validation and use the predictions on the presumed inliers as the indication of whether we believe that sample to be a mislabelled one. This procedure is shown in detail in Figure 2. We choose the amount of folds to be identical to the number of outliers to maximise the amount of confirmed outliers in the training data. The left-out outlier in the prediction data fold is used to compute the validation metrics in the results section.
To further improve the robustness of the approach and since precise performance evaluation is challenging with the limited amount of data , we use two different approaches in our ensembles. In particular, we utilise XGBoost (Chen & Guestrin, 2016) as implemented in the Python package of the same name [^6] https://xgboost.readthedocs.io/en/stable/index.html Accessed: 2024-03-08 and a balanced random forest classifier (Chen et al., 2004) as implemented in the imblearn [^7] https://imbalanced-learn.org/ Accessed: 2024-03-08 Python module. Given the invariance of both approaches to monotonous transforms, no data transforms were applied beyond the elimination of samples (i.e. sources/orbital solutions) for which any of the studied features was not available.
@@ 3.3 Feature Selection and Importance
As examples, we discarded both the parallax and the apparent magnitude of the source. These correlate positively with having confirmed substellar companions because those are the best-surveyed sources with ground-based, high-precision radial-velocity observations and generally are easier to follow-up. From an astrophysical standpoint, there is little reason for a bright source to be more likely to host a substellar companion than a faint one, at least within the parameter space surveyed here. The other confounders that we identified are indicated in Tables 5 and 6.
@@ 3.4 Ensembling and Candidate Selection
As outlined in the previous paragraphs, we use four different datasets (see Figure 1) and two different models (XGBoost and BalancedRandomForest). For each of the eight possible dataset and model combinations we compute a prediction for each source in a binary classification with value 0 to 1, where higher values mean the source is likelier to be an outlier according to the model using the scheme described in Figure 2.
These scores are not probabilities and are not calibrated, hence we cannot directly combine the output of different model and dataset combinations. Instead, we build our predictions using max-voting on the highest scoring top 50 samples per configuration. Only samples that are classified as presumed inliers are considered in those top 50. From these top 50 samples per configuration, we take those samples that occur most often across the different configurations and describe this relative occurrence rate by $\rho$ (0 if in none to 1 if in all). Note that not all samples are in all datasets and individual samples may in total occur as inliers zero (confirmed substellar companions - always outliers), two (high-probability preselected sources, orange in Figure 1), four (non-single sources) or eight (preselected sources) times. In addition to the overall relative occurrence $\rho$, which describes the rate of occurrence in the eight combinations, we observe occurrence rates $\rho_{nss}$ and $\rho_{ssc}$, which are the rates of occurrence in the combinations using large datasets (D1 & D2 containing non-single stars) and smaller ones (D3 & D4 containing only preselected sources), respectively. Details on the datasets are displayed in Figure 1.
@@ 4 Results: substellar-companion candidates
Figure 3: Density histograms of absolute magnitude as a function of colour (top panel, the inset show a zoom into the region of interest) and mass function (bottom) for all Gaia DR3 astrometric orbits. The blue contours indicate the concentration of the solutions labelled as preselected sources. Circles indicate the 10 confirmed exoplanets, squares indicate the 14 confirmed BD-companions, and crosses indicate our 22 best candidates for substellar companions discussed in the text. For reference, the Sun has $M_{G}=4.67$ and $G_{\mathrm{BP}}-G_{\mathrm{RP}}=0.82$ (Casagrande & VandenBerg, 2018), and the mass function of a $5\,M_{J}$ planet in a Jupiter-like orbit around the Sun is $1.1\cdot 10^{-7}$. These reference locations are marked with a yellow diamond. The exoplanet with the largest mass function is HD 39392 b (Wilson et al., 2016; Stevenson et al., 2023).
@@ 4.1 Validation
Ensuring the performance of our approach is a challenging task given the small amount of labelled samples. In the prediction method described in Figure 2 we obtain the predictions on the left-out fold as in a stratified K-fold validation. We use these predictions to compute performance metrics. As the inlier labels in these folds are weak labels (we expect some substellar companions in them) the metrics are probably underestimating the true performance since the models correctly "mislabel" some false inliers. Similarly, for the high probability (but not confirmed) outliers some may be mislabelled outliers, further impacting the performance measurements. Detailed results for different configurations are given in Table 3. Overall, we achieve very good results on the area under the receiver operating characteristic (AUROC) and accuracy of predictions. The achieved area under the precision recall curve (AUPRC) results are further from the optimal value of 1 but this is expected given the weak labels (in some cases the model may be right but the label wrong) and challenging task due to the significant class imbalance (less than 0.03% labelled outliers amongst all sources).
Table 3: Validation results for different datasets and models (Random Forest - RF, XGBoost - XGB) showing accuracy of prediction as well as area under the receiver operating characteristic (AUROC) and area under the precision recall curve (AUPRC). Results are averaged over the relevant runs, e.g. RF on D1, RF on D3, XGB on D1, XGB on D3 for the first column.
Datasets & Models Metric Only Confirmed Outliers D1&D3 RF+XGB High Probability Outliers D2&D4 RF+XGB All Sources D1&D2 RF+XGB Candidate Sources D3&D4 RF+XGB All Datasets D1-D4 XGB All Datasets D1-D4 RF All
@@
A secondary measure of the performance of our approach is the accuracy of the applied majority voting scheme to identify the best candidates from the eight configurations. A total of 170 unique sources show up in the top 50 outliers of the eight approaches, i.e. amongst all samples including presumed inliers and outliers. Among those 170, we find 22 of our 24 confirmed substellar companions among the 61 highest ranked samples across all configurations, each being identified in at least two separate runs. The only missing ones are the brown dwarf companions in LHS 1610 (Gaia DR3 43574131143039104, Fitzmaurice et al., 2024) and HD 89707 (Gaia DR3 3751763647996317056, Gaia Collaboration et al., 2023b). The one around LHS 1610 was likely missed due to the very-red colour of its M5-dwarf host star, which is uncommon in the confirmed substellar companions. The one around HD 89707 was barely below the detection threshold scoring highly on most models. We investigate this further in Section 4.4 on feature importance.
Table 4: Best candidates for exoplanets and brown-dwarf companions. The SpT column indicates the source’s spectral type as listed in Simbad, the solution type column indicates a short-hand for the nss_solution_type (similar to Marcussen & Albrecht 2023: OTSV is ‘OrbitalTargetedSearchValidated’, OTS is ‘OrbitalTargetedSearch’, and ASB1 is ‘AstroSpectroSB1’), $M_{2}$ is a companion-mass estimate under the dark-companion assumption and using the primary-mass estimate $M_{1}$ from binary_masses, $\rho$ is the relative occurrence introduced in Section 3.4, $\rho_{\mathrm{nss}}$ is the relative occurrence for dataset variants D1&D2 only, $\rho_{\mathrm{ssc}}$ is the relative occurrence for dataset variants D3&D4 only, the $ssc$ flag is 1 when the solution was part of ‘preselected sources ‘ or ‘high-probability preselected sources ‘ and 0 otherwise, and $M_{2,\mathrm{alt}}$ is the companion-mass range as estimated in binary_masses. The orbital solutions of * 54 Cas and BD+75 510 were retracted as false positives by the Gaia collaboration.